RC Meeting: August 29th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings


2020

March
19, 20, 21, 22, 23, 24
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

RC:M. Hattawy

Summary of Aug. 29th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3836682
  2. Issues: Extended DC HV trip in sector 5 region 1
  3. studies: x-y position on 2H01, x = -0.2mm, y -1.3mm
  4. Recorded data: 19 M on D2

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3836869
  2. Issues: the high rates; RTPC HV tripped GEM1
  3. Recorded data: 28M on D2

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3837140
  2. Studies: We are back to the beautiful running conditions. Our issue was the entrance angle of the beam to the Hall. We fixed the beam position on 2H01 then we did x and y scans on 2C24 with the BOM and the Downstream Halo counters as our observables. We had to push the beam on x to the maximum left side (-3.7 mm, 4 mm left to the nominal x value we were operating before), which is the limit of the correctors with respect to 2C21. We are running now 240nA beam current on full D2 target, getting trigger rate of 2kHz, DAQ live time of 94% and the DC occupancy is 3.5%. All the detectors' rates are going to the normal rates, which are half of what we were getting lately
  3. Recorded data: 3.3M on D2 and now switching to H2 (21:00)

Whole day summary

  1. Total recorded data today: 50 M on D2
  2. Part-II total accumulated data on D2: 1206 M events on D2
  3. Concerns: NO more concerns


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded